Hi!
I am a little lost with the problem " how many gallons of pure antifreeze must be mixed with 30 gallons of 15% antifreeze to get a mixture that is 40% antifreeze"
So far I have multiplied 30 x .15 to get 4.5 gallons of pure antifreeze but am not sure where to go after that.
Thanks!
Rachel

Answer by stanbon(75887)

how many gallons of pure antifreeze must be mixed with 30 gallons of 15% antifreeze to get a mixture that is 40% antifreeze"
-------
Equation:
alch + alch = alch
--------
0.15*30 + 1*x = 0.40(30+x)
--------
Multiply thru by 100 to get:
15*30 + 100x = 40*30 + 40x
------
60x = 25*30
x = 12.5 gallons (amt. of pure solution needed)
--------------
Cheers,
Stan H.
==============